OpenStudy (anonymous):

Dr. Garcia is twice as old as her son. Twenty Years ago, he was 4 times as old as his son was then. How old are they now?

OpenStudy (anonymous):

d = 2s d - 20 = 4(s - 20) 2s - 20 = 4s - 80 2s = 60 s = 30 Can you plug this back in and solve for d?